Identify the missing number.

22, 22, 44, 1,

13, 3, 6, ?,


1
Expert's answer
2021-04-14T14:28:33-0400

given: 22, 22, 44, 1,

13, 3, 6, ?


"\implies" "\frac{22+22}{44}=1"


"\implies" "\frac{6+3+4}{13}=1"


The missing number: "4"